Find the equation of the normal to the solid 2x^2 −y^2 +8z^2 = 11 at a point where it
intersects the line x−3 = z =(y+1)÷2.
1
Expert's answer
2020-02-18T05:59:30-0500

Solution:


"2x^2 \u2212y^2 +8z^2 = 11"

"x\u22123 = z =(y+1)\u00f72"


they intersect at points:

"(\frac{5-\sqrt{7}}{3}, \frac{-11-2\sqrt{7}}{3}, \frac{-4-\sqrt{7}}{3})"

and

"(\frac{5+\sqrt{7}}{3}, \frac{-11+2\sqrt{7}}{3}, \frac{-4+\sqrt{7}}{3})"


Normals:


"\frac{x-\frac{5-\sqrt{7}}{3}}{4\cdot\frac{5-\sqrt{7}}{3}}=\frac{y-\frac{-11-2\sqrt{7}}{3}}{-2\cdot\frac{-11-2\sqrt{7}}{3}}=\frac{z-\frac{-4-\sqrt{7}}{3}}{16\cdot\frac{-4-\sqrt{7}}{3}}"


"\frac{x-\frac{5+\sqrt{7}}{3}}{4\cdot\frac{5+\sqrt{7}}{3}}=\frac{y-\frac{-11+2\sqrt{7}}{3}}{-2\cdot\frac{-11+2\sqrt{7}}{3}}=\frac{z-\frac{-4+\sqrt{7}}{3}}{16\cdot\frac{-4+\sqrt{7}}{3}}"